# DemMamba: Alignment-free Raw Video Demoiréing with Frequency-assisted Spatio-Temporal Mamba

### v2

@@ 3.2.1 Design of SMB
Image RDNet 25.892 0.8939 0.1508 2.722 2.514
RRID 27.283 0.9029 0.1168 2.374 0.501
Video VDMoiré 27.277 0.9071 0.1044 5.836 1.057
VDMoiré* 27.747 0.9116 0.0995 5.838 1.125
DTNet 27.363 0.8963 0.1425 3.987 0.972
DTNet* 27.892 0.9055 0.1135 3.368 1.050
RawVD 28.706 0.9201 0.0904 6.585 1.247
Ours 30.004 0.9169 0.0901 2.919 0.446
3.2.2 Design of TMB
TMB, depicted in Fig. 2(b), is designed to learn inter-channel relationships, which facilitates better utilization of neighboring information and preserves temporal consistency. Serving as the core structure of TMB, the Temporal Mamba is analogous to the aforementioned SMB but employs a different scanning strategy. Specifically, we opt for forward and backward scanning [58] on the flattened temporal sequences, given that there is only one dimension on the temporal axis, as illustrated in Fig. 2(c.2). This bidirectional scanning enables effective modeling of the interrelations across the temporal domain.
To enhance temporal information interactions, we introduce a CAB to exploit the inter-channel relationship among features. As shown in Fig. 3(b), CAB consists of two standard convolutional layers with a GELU activation and a channel attention (CA) module. For computational efficiency, the number of channel in the two convolutional layers is compressed by a constant $\gamma$. For an input feature with $C$ channels, the number of channels in the output feature after the first convolutional layer is reduced to $\frac{C}{\gamma}$ and then expanded back to $C$ channels through the second layer. Subsequently, a standard CA module [54] adaptively rescales channel-wise features. Finally, the output of the TMB can be computed by the weighted summation of the output from Temporal Mamba $\mathbf{X}_{TM}$ and CAB $\mathbf{X}_{TM}$ as $\mathbf{X}_{TMB}=\mathbf{X}_{TM}+\alpha_{2}\mathbf{X}_{CAB}$, where $\alpha_{2}$ serves as the pre-defined weighting coefficient in TMB that enables the model to effectively integrate temporal dependencies while maintaining overall feature coherence.
3.3 Loss Function
@@
where $\mathbf{I}^{t}_{c}$ represents the ground-truth moiré-free image of frame $t$. We employ $L_{1}$ loss alongside the perceptual loss [17], which can closely reflect the human visual system’s perception of image quality. $\Phi_{l}(\cdot)$ denotes a set of VGG-16 layers.
4 Experiments
4.1 Experimental Setup
4.1.1 Dataset
We assess the effectiveness of our proposed DemMamba using the RawVDemoiré dataset [5], which consists of 300 clean raw source videos along with their corresponding moiréd versions in the sRGB domain. RawVDemoiré dataset is divided into 250 videos for training and 50 videos for testing. Each video comprises 60 frames at a resolution of 720p (1080×720). Four camera-screen combinations are employed for dataset collection, resulting in frames with diverse moiré patterns. In addition, we utilize the raw image demoiréing dataset TMM22 [51] to verify the generalization ability of our method. TMM22 comprises 540 paired raw and sRGB images with ground truth for training and 408 pairs for testing, covering various recaptured scenes such as natural images, webpages, and documents. To facilitate the training and comparison process, patches sized 256$\times$256 and 512$\times$512 are cropped for the training and testing sets, respectively.
4.1.2 Training Details
DemMamba is trained using the AdamW optimizer, with $\beta_{1}$ set to 0.9 and $\beta_{2}$ set to 0.999. We employ a multistep learning rate schedule, with the learning rate initialized at $4\times 10^{-4}$. DemMamba is trained for 50 epochs using a batch size of 8 on 2 NVIDIA Tesla H800 GPUs. For the structure in DemMamba, the numbers of FASTMG, SMB, and TMB are all set to 4. The weighting coefficient in SMB($\alpha_{1}$) and TMB($\alpha_{2}$) are both set to 0.5. In CAB, we set the squeeze factor $\gamma$ between the two convolutional layers to 3.
@@ 4.1.2 Training Details
4.2 Quantitative Results
We compare our approach with two raw image demoiréing methods, RDNet [51] and RRID [48], along with several video demoiréing methods: VDmoiré [6], DTNet [47] and RawVD [5]. Since VDmoiré and DTNet are originally designed for video demoiréing in the sRGB domain, we have retrained them with raw inputs, adjusting the input channel number and inserting upsampling layers to adapt them to raw data. The revised versions are marked with an $\ast$. To evaluate the performance of our proposed DemMamba, we adopt the following three standard metrics to assess pixel-wise accuracy and the perceptual quality: PSNR, SSIM [43], and LPIPS [53]. Additionally, we assess model complexity by analyzing the number of parameters and inference time. To ensure a fair comparison, all models are fine-tuned using the default settings provided in their respective papers. We select the better results from the pretrained and retrained models for comparison, providing an advantage over competing methods.
Table 1 presents the quantitative comparisons and computational complexity analysis on the RawVDmoiré dataset [5]. Our proposed DemMamba outperforms the second-best method, RawVD, by 1.3 dB in PSNR and 0.0003 in LPIPS. For SSIM, our method achieves the second-best result with a score of 0.9169, surpassing most of the compared methods. Regarding computational cost, our approach achieves effective demoiréing with the inference time of merely 0.446 seconds. Compared to previous methods, DemMamba demonstrates superiority with an acceptable parameter count of 2.919M, further highlighting the advantages of our approach.
@@ 4.2 Quantitative Results
DMCNN 23.54 0.885 0.154
@@
WDNet 22.33 0.802 0.166
[…]
ESDNet 26.77 0.927 0.089
[…]
RDNet 26.16 0.921 0.091
RRID 27.24 0.929 0.098
RawVD 27.26 0.935 0.075
Ours 28.14 0.936 0.067
To verify the generalization performance of our method, we conduct additional experiments on the raw image demoiréing dataset TMM22 [51]. To adapt to the TMM22 dataset, we input the same three frames into our model. In Table 2, we compare our method with several image demoiréing methods, including DMCNN [37], WDNet [24], and ESDNet [50] (along with their corresponding raw input versions, denoted with $\ast$), as well as raw image demoiréing methods RDNet [51] and RRID [48], and the raw video demoiréing method RawVD [5]. Although DemMamba is initially designed for raw video demoiréing, it achieves the best performance on TMM22 dataset with a PSNR of 28.14 dB and an SSIM of 0.936. Regarding perceptual quality, DemMamba surpasses all the other methods with an LPIPS score of 0.067. The qualitative results on the TMM22 dataset further demonstrate the generalization capability of our method.
4.3 Qualitative Results
We present visual comparisons between our approach and existing methods in Fig. 4. The results demonstrate DemMamba’s ability in removing moiré patterns and correcting color deviations. In the first scenario, noticeable vertical moiré stripes appear in the light background regions. Unfortunately, the images restored by RDNet, RRID, VDmoiré and DTNet still exhibit remaining moiré patterns. Although RawVD can eliminate most of the moiré patterns, the resulting image tends to have a slight grayish bias. In contrast, DemMamba effectively removes the moiré patterns, preserves the original image details, and restores accurate colors. For the second scenario, which includes dark backgrounds with limited spatial details for model learning, our method still demonstrates stable demoiréing performance. DemMamba maintains higher texture accuracy, color fidelity, and structural coherence compared to other methods, underscoring its robustness in handling challenging low-light conditions.
Additionally, we present the qualitative comparison on raw image demoiréing dataset TMM22 in Fig. 5. In the first scene, which features a blue umbrella with moiré patterns, previous methods tend to leave residual moiré patterns in their restored images. In the final scene, we present a case of webpages affected by moiré patterns. Although some methods partially remove moiré patterns, they struggle to restore the original colors accurately. DemMamba’s ability to model long-range dependencies enables a more accurate restoration of colors, resulting in enhanced visual quality. Additional image and video comparisons are provided in the supplementary file.
Table 3: Ablation study for investigating the components in our proposed DemMamba.
[…]
MambaIR 29.01 0.9119
w/o AFB 29.74 0.9161
w/o CAB 29.92 0.9167
all SMB 29.65 0.9156
all TMB 28.65 0.9118
DemMamba 30.00 0.9169
4.4 Ablation Study
Table 3 shows an assessment of the effectiveness of our proposed DemMamba through ablation experiments with various combinations of the foundational elements. Additionally, we introduce MambaIR [12] as a key baseline for comparison. Firstly, we validate the performance of MambaIR, which serves as an Mamba-based image restoration baseline. MambaIR achieves a PSNR of 29.01 dB in the raw video demoiréing task, demonstrating Mamba’s capability to manage long-term dependencies. However, the design of MambaIR limits its capacity in capturing temporal features, which is critical for video demoiréing. DemMamba employs alternating SMB and TMB to effectively capture intra- and inter-frame dependencies in raw videos affected by moiré patterns. To validate the effectiveness of this specific design, we replace one block type with the other. For instance, substituting all SMBs with TMBs results in a drastic PSNR decrease of approximately 1.4 dB. In addition, removing AFM and CAB respectively from the SMB and TMB leads to slight decreases of 0.26 dB and 0.08 dB in PSNR, underscoring the importance of the two modules. Incorporating all the designs, our complete model surpasses MambaIR by nearly 1 dB in PSNR, thereby affirming the efficacy of our frequency-assisted spatio-temporal Mamba.
5 Conclusion
@@
We propose the innovative DemMamba network, an alignment-free raw video demoiréing method enhanced by frequency-assisted spatio-temporal Mamba. DemMamba adapts the Mamba framework to video restoration, strategically introducing SMB and TMB to model the inter- and intra-relationships, mitigating moiré patterns while ensuring temporal consistency. SMB captures global correlations within the spatial domain. Considering the frequency nature of moiré patterns, AFM is integrated as a learnable compressor, automatically adapting to and reducing specific moiré frequencies. TMB enhances the utilization of adjacent data to maintain temporal stability, supplemented by a CAB that strengthens temporal information integration. Extensive experiments demonstrate that our proposed DemMamba outperforms state-of-the-art methods both qualitatively and quantitatively.

### v3

@@ Introduction
Image RDNet 25.892 0.8939 0.1508 553.50 0.9130 2.722 2.514
RRID 27.283 0.9029 0.1168 311.85 0.9306 2.374 0.501
Video VDMoiré 27.277 0.9071 0.1044 405.22 0.9158 5.836 1.057
[…]
DTNet 27.363 0.8963 0.1425 573.34 0.9174 3.987 0.972
[…]
RawVD 28.706 0.9201 0.0904 355.75 0.9246 6.585 1.247
Ours 30.069 0.9192 0.1115 289.48 0.9350 2.941 0.592
Ours-L 30.317 0.9195 0.0900 276.29 0.9361 4.382 0.686
Table 1: Comparison with state-of-the-art image and video demoiréing methods for raw video demoiréing in terms of PSNR, SSIM, LPIPS, FVD, FSIM and computing complexity. The best result is highlighted in bold and the second best is underlined.
Design of TMB
TMB, depicted in Fig. 3(b), is designed to learn inter-frame relationships that enhance the utilization of neighboring information, thereby alleviating moire patterns and preserving temporal consistency throughout the video sequence. Serving as the core structure of TMB, the Temporal Mamba is analogous to the aforementioned Spatial Mamba but employs a different scanning strategy. Specifically, we opt for forward and backward scanning (Zhu et al. 2024a) on the flattened temporal sequences, given that the temporal axis is one-dimensional, as illustrated in Fig. 3(c.2). This bidirectional scanning effective models the interrelations across the temporal domain.
To enhance temporal information interactions, we introduce a CAB to exploit the inter-channel relationship among features. As shown in Fig. 4(b), CAB consists of a set of convolutional layers with different kernel size ($k=1,3,5$) and a channel attention (CA) module. We utilize convolutional kernels of varying sizes to capture temporal dependencies across diverse temporal length. This parallel architecture enables the network to learn the most effective kernel configurations, thereby enhancing its ability to model complex temporal patterns. Subsequently, a standard CA module (Zhang et al. 2018b) adaptively rescales channel-wise features. Finally, the output of the TMB can be computed by the weighted summation of the output from Temporal Mamba $\mathbf{X}_{TM}$ and $\mathbf{X}_{CAB}$ as $\mathbf{X}_{TMB}=\mathbf{X}_{TM}+\alpha_{2}\mathbf{X}_{CAB}$, where $\alpha_{2}$ serves as the pre-defined weighting coefficient in TMB that enables the model to effectively integrate temporal dependencies while maintaining overall feature coherence.
Loss Function
@@
where $\mathbf{I}^{t}_{c}$ represents the ground-truth moiré-free image of frame $t$. We employ $L_{1}$ loss alongside the perceptual loss (Johnson, Alahi, and Fei-Fei 2016), which can closely reflect the human visual system’s perception of image quality. $\Phi_{l}(\cdot)$ denotes a set of VGG-16 layers.
Experiments
Experimental Setup
Dataset. We assess the effectiveness of our proposed DemMamba using the RawVDemoiré dataset (Cheng, Liu, and Yang 2024), which consists of 300 clean raw source videos along with their corresponding moiréd versions in the sRGB domain. RawVDemoiré dataset is divided into 250 videos for training and 50 videos for testing. Each video comprises 60 frames at a resolution of 720p (1080×720). Four camera-screen combinations are employed for dataset collection, resulting in frames with diverse moiré patterns. In addition, we utilize the raw image demoiréing dataset TMM22 (Yue et al. 2022) to verify the generalization ability of our method. TMM22 comprises 540 paired raw and sRGB images with ground truth for training and 408 pairs for testing, covering various recaptured scenes such as natural images, webpages, and documents. To facilitate the training and comparison process, patches sized $\times256$ and $\times512$ are cropped for the training and testing sets, respectively.
Training Details. DemMamba is trained using the AdamW optimizer, with $\beta_{1}$ set to 0.9 and $\beta_{2}$ set to 0.999. We employ a multistep learning rate schedule, with the learning rate initialized at $4\times 10^{-4}$. DemMamba is trained for 100 epochs using a batch size of 8 on 4 NVIDIA Tesla A800 GPUs. For the structure in DemMamba, the numbers of FASTMG, SMB, and TMB are all set to 4. The weighting coefficient in SMB($\alpha_{1}$) and TMB($\alpha_{2}$) are both set to 0.5.
Quantitative Results
Frame Level Comparison. We compare our approach with two raw image demoiréing methods, RDNet (Yue et al. 2022) and RRID (Xu et al. 2024a), along with several video demoiréing methods: VDmoiré (Dai et al. 2022), DTNet (Xu et al. 2024b) and RawVD (Cheng, Liu, and Yang 2024). Since VDmoiré and DTNet are originally designed for video demoiréing in the sRGB domain, we have retrained them with raw inputs, adjusting the input channel number and inserting upsampling layers to adapt them to raw data. The revised versions are marked with an $\ast$. To evaluate the performance of our proposed DemMamba, we adopt the following three standard metrics to assess pixel-wise accuracy and the perceptual quality: PSNR, SSIM (Wang et al. 2004), and LPIPS (Zhang et al. 2018a). Additionally, we assess model complexity by analyzing the number of parameters and inference time. To ensure a fair comparison, all models are fine-tuned using the default settings provided in their respective papers. We select the better results from the pretrained and retrained models for comparison, providing an advantage over competing methods.
@@
Table 1 presents the quantitative comparison and computational complexity analysis on the RawVDmoiré dataset. Our proposed method, DemMamba, surpasses the second-best approach, RawVD, achieving a PSNR of 30.069 dB with an inference time of only 0.686 seconds, demonstrating both efficiency and effectiveness in demoiréing. Furthermore, DemMamba maintains this performance with a modest parameter count of 2.941M. Additionally, we provide a larger version of our model, denoted as Ours-L, achieves an additional 1.6 dB improvement in PSNR over RawVD. Improvements are also observed in SSIM and LPIPS metrics, further substantiating the superiority of our approach.
@@
[…]
DMCNN 23.54 0.885 0.154 0.052
[…]
WDNet 22.33 0.802 0.166 0.284
[…]
ESDNet 26.77 0.927 0.089 0.115
[…]
RDNet 26.16 0.921 0.091 1.094
RRID 27.24 0.929 0.098 0.326
RawVD 27.26 0.935 0.075 0.182
Ours 28.14 0.936 0.067 0.159
@@
To verify the generalization of our method, we conduct additional experiments on the raw image demoiréing dataset TMM22 (Yue et al. 2022). To adapt to the TMM22 dataset, we input the same three frames into our model. In Table 2, we compare our method with several image demoiréing methods, including DMCNN (Sun, Yu, and Wang 2018), WDNet (Liu et al. 2020a), and ESDNet (Yu et al. 2022) (along with their corresponding raw input versions, denoted with $\ast$), as well as raw image demoiréing methods RDNet and RRID, and the raw video demoiréing method RawVD. Although DemMamba is initially designed for raw video demoiréing, it achieves the best performance on TMM22 dataset with a PSNR of 28.14 dB and an SSIM of 0.936. Regarding perceptual quality, DemMamba surpasses all the other methods with an LPIPS score of 0.067. The qualitative results on the TMM22 dataset demonstrate the generalization capability of our method.
Video-Level Comparison. We utilize FVD  (Unterthiner et al. 2018) and FSIM  (Zhang et al. 2011) to assess the quality of video outputs. FVD, which extends the Fréchet Inception Distance concept, evaluates the temporal coherence of videos. FSIM is an image quality assessment index that evaluates changes in image quality by comparing the similarity of low-level content features, such as gradients and phase consistency. As indicated in Table 1, DemMamba outperforms existing approaches in both metrics, demonstrating that our outputs more closely resemble the target distribution of entire video sequences and better preserve per-pixel and structural visual information. Additionally, we have included a video file in the supplementary material to provide a visual comparison of video quality.
Qualitative Results
Visual comparisons between the proposed method and existing methods are presented in Fig. 5. The results demonstrate DemMamba’s effectiveness in removing moiré patterns and correcting color deviations. Notably, vertical moiré stripes are visible in the background regions. Images restored by RDNet, RRID, VDmoiré, and DTNet still contain residual moiré artifacts. Although RawVD can eliminate most moiré patterns, the restored images exhibit a slight grayish bias. In contrast, DemMamba effectively removes moiré patterns while accurately restoring colors.
Additionally, we present the qualitative comparison on raw image demoiréing dataset TMM22 in Fig. 6. In the first scene, which features a blue umbrella with moiré patterns, previous methods tend to leave residual moiré patterns in their restored images. In the second scene, we present a case of webpages affected by moiré patterns. Although some methods partially remove moiré patterns, they struggle to restore the original colors accurately. DemMamba’s ability to model long-range dependencies enables a more accurate restoration of colors, resulting in enhanced visual quality.
Ablation Study
Different design choices of FASTMG. Table 3 shows an assessment of DemMamba through ablation experiments with various combinations of the foundational elements. Replacing TMB with SMB results in a PSNR of 30.15 dB and an SSIM of 0.9187, whereas the opposite substitution (replace SMB with TMB) further degrades performance to 29.92 dB and 0.9179, confirming that neither block alone can fully capture both spatial and temporal correlations. Furthermore, removing AFB and CAB respectively leads to slight performance decreases, underscoring the importance of the two modules in moire pattern removal and temporal information utilization. In Fig. 7, we also provide visual representations to further elucidate the effectiveness of the proposed modules in removing moiré patterns.
@@
replace TMB with SMB 30.15 0.9187
replace SMB with TMB 29.92 0.9179
remove AFB 30.22 0.9192
remove CAB 30.18 0.9190
DemMamba (Ours) 30.32 0.9195
Table 3: Different design choices of FASTMG.
Figure 7: Different design choices of FASTMG.
Different scan modes in SMB. Table 4 reports the performance of the SMB when employing varying numbers of scan modes. Increasing the number of scan directions from one (1-d) to eight (8-d) yields improvements of 0.57 dB in PSNR and 0.0026 in SSIM, with acceptable increases in parameter count and runtime. Moreover, Fig. 8 visually demonstrates that increasing the number of scan directions progressively reduces moiré artifacts across multiple orientations, further substantiating the effectiveness of the proposed multi-directional scanning strategy.
@@
1-d 29.75 0.9169 3.797 0.204
2-d 29.98 0.9182 3.881 0.221
4-d 30.24 0.9191 4.048 0.454
8-d 30.32 0.9195 4.382 0.686
Table 4: Different scan modes in SMB.
Figure 8: Different scan modes in SMB.
Conclusion
@@
We propose DemMamba, an alignment-free raw video demoiréing method enhanced by frequency-assisted spatio-temporal Mamba. DemMamba introduces SMB and TMB to model the inter- and intra-relationships, mitigating moiré patterns while ensuring temporal consistency. The SMB employs a multi-directional scanning mechanism coupled with a learnable frequency compressor to differentiate interference patterns across various orientations and frequencies. TMB enhances the utilization of adjacent data to maintain temporal stability, supplemented by a CAB that strengthens temporal information integration. Extensive experiments demonstrate that DemMamba outperforms state-of-the-art methods both qualitatively and quantitatively.